What is the number which when multiplied by 2 and added to 8 gives the same result as when it is divided by 2 and has 32 added to it ?

Answer by fcabanski(1391)

Call the number x.


2x+8=x/2 +32


2x - x/2 = 32-8


1.5x = 24


x=16